# One-to-one correspondence reconstruction at the electron-positron Higgs factory

### v1

@@ Abstract
We propose one-to-one correspondence reconstruction for electron-positron Higgs factories. For each visible particle, one-to-one correspondence aims to associate relevant detector hits with only one reconstructed particle and accurately identify its species. To achieve this goal, we develop a novel detector concept featuring 5-dimensional calorimetry that provides spatial, energy, and time measurements for each hit, and a reconstruction framework that combines state-of-the-art particle flow and artificial intelligence algorithms. In the benchmark process of Higgs to di-jets, over 90% of visible energy can be successfully mapped into well-reconstructed particles that not only maintain a one-to-one correspondence relationship but also associate with the correct combination of cluster and track, improving the invariant mass resolution of hadronically decayed Higgs bosons by 25%. Performing simultaneous identification on these well-reconstructed particles, we observe efficiencies of 97% to nearly 100% for charged particles ($e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$) and photons ($\gamma$), and 75% to 80% for neutral hadrons ($K_{L}^{0}$, $n$, $\bar{n}$). For physics measurements of Higgs to invisible and exotic decays, golden channels to probe new physics, one-to-one correspondence could enhance discovery power by 10% to up to a factor of two. This study demonstrates the necessity and feasibility of one-to-one correspondence reconstruction at electron-positron Higgs factories.
In collider experiments, high-energy particles impact each other and generate multiple final state particles, bringing the information of underlying physics laws. These particles interact with the detector, excite detector sensors, and generate electronic signals that are recorded into data, typically in the form of detector hits. Correspondingly, data processing in collider experiments consists of two steps: interpreting detector hits into physics objects, especially final state particles, and performing physics measurements using these physics objects. The first step is called reconstruction, which can be regarded as the inverse process of detector hits generation.
@@
Depending on the multiplicity of visible particles and the detector configuration, achieving 1-1 correspondence could be truly challenging. For proton-proton and heavy ion collisions at the LHC [1] or future hadron colliders, each collision can generate $\mathcal{O}(10^{2}\text{--}10^{4})$ visible particles, making 1-1 correspondence extremely difficult, especially in the detector forward region. In contrast, at electron-positron colliders with center-of-mass energies ($\sqrt{s}$) around 10 GeV or lower, such as the BelleII [2, 3] and BESIII [4, 5] experiments, the particle multiplicity is typically below 10, and their detectors usually emphasize on the PID performance, achieving 1-1 correspondence could be much easier. Our study focuses on the electron-positron Higgs factory [6, 7, 8, 9, 10] which is regarded as the highest-priority future collider as it offers excellent opportunity to discover new physics [11, 12], especially via precision measurements of the Higgs boson. It operates at center-of-mass energies ranging from 91.2 GeV to several TeV, typically producing $\mathcal{O}(100)$ visible particles distributed in narrow jets in hadronic events.
The concept of 1-1 correspondence evolves from the Particle Flow Algorithm (PFA) [13, 14, 15, 16], with the key idea to trace each individual particle, classify and associate corresponding detector hits, and measure the particle 4-momentum using the most suitable sub-detector system. Originating from the ALEPH [16] experiment, the PFA has become the guiding principle for multiple conceptual detector designs at electron-positron Higgs factories [9, 10, 17] and detector upgrade projects at HL-LHC [18]. Benefit from the rapid development of artificial intelligence, many machine learning algorithms have been implemented to enhance the PFA performance [19, 20, 21]. The current PFA performance is primarily limited by the confusion [22], which includes the effects of fake particles, failures in track-cluster matching, particle loss due to shower overlap, etc. These confusion effects violate the 1-1 correspondence relation. Therefore, the goal of 1-1 correspondence reconstruction is to approach confusion-free particle flow reconstruction and to identify the species of all visible particles.
@@
Figure 1: AURORA detector geometry and an event display of $\nu\bar{\nu}H,H\to gg$ ($\sqrt{s}$ = 240 GeV) event.
@@
The performance of PFA can be quantified using the Boson Mass Resolution (BMR) [9, 23], which represents the relative mass resolution of massive bosons (such as Higgs, $Z$, and $W$ bosons) decaying into hadronic final states. At the electron-positron Higgs factory, a BMR smaller than 4% is required to separate the $q\bar{q}H$ signal from $q\bar{q}Z$ background using the recoil mass method [9, 24]. A smaller BMR is beneficial for all physics measurements with hadronic final states, especially for flavor physics measurements and new physics searches since they rely on accurate measurements of missing energy and momentum. Taking reference to the Circular Electron-Positron Collider (CEPC) Conceptual Design Report (CDR) [9], its baseline detector design achieves a BMR of 3.7% [25, 23], meeting the requirements for Higgs measurements. In terms of PID, the CEPC baseline detector demonstrates a typical efficiency of 99.5% and a misidentification rate around 1% for isolated lepton identification [26, 27]. It can also distinguish charged hadrons ($\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$) using techniques like ${\rm d}E/{\rm d}x$ and time-of-flight (TOF) measurements [28, 29].
To pursue 1-1 correspondence reconstruction, we develop the AURORA (ApparatUs for RecOnstRuction with Advanced algorithm) detector concept and PROOF (Particle Reconstruction with One-to-One correspondence at Higgs Factory) reconstruction framework, based on the CEPC CDR baseline design [9]. AURORA features a high-granularity 5-dimensional (5D) calorimetry as its core, with a unified time resolution of 100 ps for each calorimeter hit. The left panel of Fig. 1 shows a cut-away view of AURORA, with more details provided in Methods. Combining the CEPC baseline PFA Arbor [14] and Particle Transformer (ParT) [30], PROOF performs a universal identification of all reconstructed particles, simultaneously distinguishing different confusion types and identifying particle species. ParT is a particle cloud- [31] and Transformer-based [32] machine learning model adapted for high-energy physics experiments, which has been applied in jet flavor tagging at the LHC [33, 34, 35] and jet origin identification (JOI) at the electron-positron Higgs factory [36]. Following the convention of CEPC performance studies [9, 25, 37], we simulate 1 million $\nu\bar{\nu}H,H\to gg$ ($\sqrt{s}$ = 240 GeV) events with AURORA detector and reconstruct these events with Arbor. A reconstructed $\nu\bar{\nu}H,H\to gg$ event is displayed in the right panel of Fig. 1, and more details on simulation and reconstruction are described in Methods.
@@
Using truth-level information in the Monte Carlo (MC) simulation, we establish a mapping between visible particles and reconstructed ones, as illustrated in the left panel of Fig. 2. The mapping consists of three main categories:
  • 1-1 correspondence (solid arrows in Fig. 2): one visible particle is mapped to only one reconstructed particle. This category consists of three types, represented in shades of green in Fig. 2. One type is well-reconstructed, where a cluster is a must for neutral particles, and both track and cluster are required for charged ones. The other two types contain charged particles that generate only track or cluster, mainly caused by the limitations of detector acceptance (e.g. in the forward region) and efficiency (e.g. dead zone or energy/momentum measurement thresholds).
  • Fake particles (in orange and yellow in Fig. 2): occur when one visible particle is mapped to multiple reconstructed particles. Due to detector inefficiency and particle interactions with upstream materials, the calorimeter hits generated by a single particle may be grouped into multiple clusters. For a charged particle, its track could be matched to only part of or even none of these clusters, while the remaining clusters are reconstructed as nearby extra neutral particles. These neutral particles that originate from charged particle shower fragments and un-associated clusters, represented in orange in Fig. 2, can cause a double-counting in the total reconstructed energy. For simplicity, these neutral particles are called fake particles in this study.
According to this mapping with truth information, we label reconstructed particles into 15 types, including 10 types of visible particles ($e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$, $\gamma$, $K_{L}^{0}$, $n$, $\bar{n}$, and others including $K_{S}^{0}$ and $\Lambda$ that not yet decay within the tracker volume) and 5 types of confusion (such as different kinds of fake particles, charged particle without track, etc). Each reconstructed particle is characterized by 55 observable variables as inputs to ParT. More details can be found in Methods.
The quantitative analysis [38] shows that fake particles are the leading contribution to the BMR. In our simulation, on average, a $\nu\bar{\nu}H,H\to gg$ event has 80 visible particles with a total energy of 135 GeV at the truth level, while the reconstructed fake particles lead to a double-counted energy of 6 GeV, severely degrading the BMR. PROOF can identify the fake particles with a typical efficiency of 77% and a purity of 97.5%, consequently suppressing the average double-counted energy from fake particles to 0.7 GeV (by nearly one order of magnitude) at the minor cost of mis-excluding visible particles with a total energy on the order of 0.5 GeV. Consequently, PROOF achieves a BMR of 2.75% (see the left panel of Fig. 3), where the effects of fake and mis-excluded particles get balanced. Compared to the CEPC CDR [39], the BMR is improved by 25%, where 10% is from the detector geometry optimization to AURORA [40], and 15% is from the fake particle suppression. The contributions from the three categories of visible-reconstructed particle mapping to the BMR are not proportional to their corresponding energy fractions. The contribution from the 1-1 correspondence category primarily stems from the detector resolution, which is orders of magnitude lower compared to the total energy of particles within this category. Meanwhile, the contributions from the fake and lost categories are comparable to their corresponding total energies. We found that the 1-1 correspondence category contributes to half of the BMR, while the fake and lost categories each contribute a quarter, combined quadratically.
@@
Referring to the truth level energy of visible particles, we calculate the fraction of total visible energy corresponding to different categories of the mapping. Ignoring the remaining fake particles [^2] The remaining fake particles could violate the 1-1 correspondence relationship, as one visible particle could be mapped to one leading reconstructed particle together with one or even several fake particles. However, as the energy of the fake particle is usually much lower than the leading one, the fake particle only has a tiny impact on the reconstruction and identification of the original visible particle. Meanwhile, the total energy of remaining fake particles only contributes to a sub-percentage level of total visible energy. Therefore, we ignore these remaining fake particles in our study., we observe that over 90% of the total visible energy is mapped to the well-reconstructed particles, 4% to charged particles without clusters (with over half of this due to calorimeter acceptance/inefficiency), 2.5% to charged particles without tracks (mostly in the forward region), and 2.6% of the energy is lost due to limited detector acceptance, shower merging, and mis-excluded genuine particles, as illustrated in the pie chart in Fig. 2. The detector acceptance dominates the confusion of charged particles without clusters or tracks. Summing the energy fraction of well-reconstructed particles and confusion type purely originating from detector acceptance, we consider that nearly 95% of the total visible energy maintains the 1-1 correspondence relationship.
@@
It should be remarked that many simplifications are made to facilitate the analysis. To truly realize 1-1 correspondence reconstruction, dedicated studies are required in the future. For instance, quantifying sub-detector performance specifications—such as the time (TOF) resolution of calorimeters, ${\rm d}E/{\rm d}x$ resolution of gaseous trackers, and detector geometry optimization—is essential. The reconstruction performance of the primary vertex is also relevant for TOF measurements. The impact of beam-induced background needs to be quantified. Modeling and verifying detailed detector responses are crucial for controlling systematic uncertainties of MC simulations. Secondary particles, treated in the same manner as primary particles in this study, can significantly affect reconstruction performance and physics measurements. For example, electron-positron pairs from photon conversion and muons from hadron decays could contribute to the background in measurements that depend on (semi-)leptonic decays. Therefore, these secondary particles require dedicated identification methods, which could be developed within the future PROOF framework.
Prospects
Our study demonstrates the feasibility of 1-1 correspondence reconstruction at the electron-positron Higgs factory, it can significantly improve the BMR performance, and simultaneously identify 10 different kinds of visible particles. Meanwhile, the performance bottleneck of 1-1 correspondence in our current setup is identified to be the detector acceptance and neutral hadron identification, which could be ameliorated via dedicated algorithms and detector optimization. This section discusses the prospects for high-level reconstruction and physics measurements with an ideal 1-1 correspondence reconstruction.

### v4

@@ Abstract
We propose one-to-one correspondence reconstruction for electron-positron Higgs factories. For each visible particle, one-to-one correspondence aims to associate relevant detector hits with only one reconstructed particle and accurately identify its species. To achieve this goal, we develop a novel detector concept featuring 5-dimensional calorimetry that provides spatial, energy, and time measurements for each hit, and a reconstruction framework that combines state-of-the-art particle flow and machine learning algorithms. In the benchmark process of Higgs to di-jets, over 91% of visible energy can be successfully mapped into well-reconstructed particles that not only maintain a one-to-one correspondence relationship but also associate with the correct combination of cluster and track, improving the invariant mass resolution of hadronically decayed Higgs bosons by 25%. Performing simultaneous identification on these well-reconstructed particles, we observe efficiencies of 97% to nearly 100% for charged particles ($e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$) and photons ($\gamma$), and 75% to 80% for neutral hadrons ($K_{L}^{0}$, $n$, $\bar{n}$). For physics measurements of Higgs to invisible and exotic decays, golden channels to probe new physics, one-to-one correspondence could enhance discovery power by 10% to up to a factor of two. This study demonstrates the necessity and feasibility of one-to-one correspondence reconstruction at electron-positron Higgs factories.
keywords:
Reconstruction , Particle Flow , Machine Learning , Higgs Factory
journal: Computer Physics Communications
[…]
1 Introduction
In collider experiments, high-energy particles collide and generate multiple final state particles, bringing information about underlying physics laws. These particles interact with the detector, excite detector sensors, and generate electronic signals that are recorded into data, typically in the form of detector hits. Correspondingly, data processing in collider experiments consists of two main steps: interpreting detector hits into physics objects, particularly final state particles, and performing physics measurements using these physics objects. The first step is called reconstruction, which can be regarded as the inverse process of detector hits generation.
@@ 1 Introduction
Figure 1: AURORA detector geometry and an event display of a reconstructed $e^{+}e^{-}\to Z(\to\nu\bar{\nu})H(\to gg)$ ($\sqrt{s}$ = 240 GeV) event.
Depending on the multiplicity of visible particles and the detector configuration, achieving 1-1 correspondence could be truly challenging. For proton-proton and heavy ion collisions at the Large Hadron Collider (LHC) [1] or future hadron colliders, each collision can generate $\mathcal{O}(10^{2}\text{--}10^{4})$ visible particles, making 1-1 correspondence extremely difficult, especially in the detector forward region. In contrast, at electron-positron colliders with center-of-mass energies ($\sqrt{s}$) around 10 GeV or lower, such as the BelleII [2, 3] and BESIII [4, 5] experiments, the particle multiplicity is typically below 10, and their detectors usually emphasize on the PID performance, achieving 1-1 correspondence could be much easier. Our study focuses on the electron-positron Higgs factory [6, 7, 8, 9, 10], which is regarded as the highest-priority future collider as it offers excellent opportunity to discover new physics [11, 12], especially via precision measurements of the Higgs boson. It operates at center-of-mass energies ranging from 91.2 GeV to several TeV, typically producing $\mathcal{O}(100)$ visible particles distributed in narrow jets in hadronic events.
The concept of 1-1 correspondence evolves from the Particle Flow Algorithm (PFA) [13, 14, 15, 16, 17, 18], with the key idea to trace each individual particle, classify and associate corresponding detector hits, and measure the particle 4-momentum using the most suitable sub-detector system. Originating from the ALEPH [16] experiment, the PFA has become the guiding principle for multiple conceptual detector designs at electron-positron Higgs factories [9, 10, 19] and detector upgrade projects at HL-LHC [20]. Benefiting from the rapid development of artificial intelligence, many machine learning algorithms have been implemented to enhance PFA performance [21, 22, 23]. The current PFA performance at electron-positron Higgs factories is primarily limited by confusion [14, 24], which includes the effects of fake particles, failures in track-cluster matching, particle loss due to shower overlap, etc. These confusion effects violate the 1-1 correspondence relationship. Therefore, the goal of 1-1 correspondence reconstruction is to approach confusion-free particle flow reconstruction and to identify the species of all visible particles.
@@
The performance of PFA can be quantified using the Boson Mass Resolution (BMR) [9, 25], which represents the relative mass resolution of massive bosons (such as Higgs, $Z$, and $W$ bosons) decaying into hadronic final states. At the electron-positron Higgs factory, a BMR smaller than 4% is required to separate the $q\bar{q}H$ signal from $q\bar{q}Z$ background using the recoil mass method [9, 26]. A smaller BMR is beneficial for all physics measurements with hadronic final states, especially for flavor physics measurements and new physics searches since they rely on accurate measurements of missing energy and momentum. Referring to the Circular Electron-Positron Collider (CEPC) Conceptual Design Report (CDR) [9], its baseline detector design achieves a BMR of 3.7% [27, 25], meeting the requirements for Higgs measurements. In terms of PID, the CEPC baseline detector demonstrates a typical efficiency of 99.5% and a misidentification rate around 1% for isolated lepton identification [28, 29]. It can also distinguish charged hadrons ($\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$) using techniques such as ${\rm d}E/{\rm d}x$ and time-of-flight (TOF) measurements [30, 31].
@@
To pursue 1-1 correspondence reconstruction, we develop the AURORA (ApparatUs for RecOnstRuction with Advanced algorithm) detector concept and the PROOF (Particle Reconstruction with One-to-One correspondence at Higgs Factory) reconstruction framework, based on the CEPC CDR baseline design [9]. AURORA features a high-granularity 5-dimensional (5D) calorimetry as its core, with a unified time resolution of 100 ps for each calorimeter hit. The left panel of Fig. 1 shows a cut-away view of AURORA, with more details provided in A. Combining the CEPC baseline PFA Arbor [14] and Particle Transformer (ParT) [32], PROOF performs a universal identification of all reconstructed particles, simultaneously distinguishing different confusion types and identifying particle species. ParT is a particle cloud- [33] and Transformer-based [34] machine learning model adapted for high-energy physics experiments, which has been applied in jet flavor tagging at the LHC [35, 36, 37] and jet origin identification (JOI) at the electron-positron Higgs factory [38]. Following the convention of CEPC performance studies [9, 27, 39], we simulate 1 million $e^{+}e^{-}\to Z(\to\nu\bar{\nu})H(\to gg)$ ($\sqrt{s}$ = 240 GeV) events (abbreviated as $ZH\to\nu\bar{\nu}gg$ in the following text) with AURORA detector and reconstruct these events with Arbor. A reconstructed $ZH\to\nu\bar{\nu}gg$ event is displayed in the right panel of Fig. 1, and more details on simulation and reconstruction are described in A.
2 Methods and results
2.1 Particle mapping
@@
Using truth-level information in the Monte Carlo (MC) simulation, we establish a mapping between simulated visible particles and reconstructed ones, as illustrated in the left panel of Fig. 2. The mapping consists of three main categories:
  • 1.
    1-1 correspondence (solid arrows in Fig. 2): one visible particle is mapped to only one reconstructed particle. This category consists of three types, represented in shades of green in Fig. 2. One type is well-reconstructed, where a cluster is a must for neutral particles, and both track and cluster are required for charged ones. The other two types correspond to charged particles that generate only track or cluster, mainly caused by the limitations of detector acceptance (e.g. in the forward region) and efficiency (e.g. dead zone or energy/momentum measurement thresholds).
  • 2.
    Fake particles (in orange and yellow in Fig. 2): occur when one visible particle is mapped to multiple reconstructed particles. Due to detector inefficiency and particle interactions with upstream materials, the calorimeter hits generated by a single particle may be grouped into multiple clusters. For a charged particle, its track could be matched to only part of or even none of these clusters, while the remaining clusters are reconstructed as nearby extra neutral particles. These neutral particles that originate from charged particle shower fragments and un-associated clusters, represented in orange in Fig. 2, can cause double-counting in the total reconstructed energy. For simplicity, these neutral particles are called fake particles in this study.
  • 3.
According to this mapping with truth information, we label reconstructed particles into 15 types, including 10 types of visible particles ($e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$, $\gamma$, $K_{L}^{0}$, $n$, $\bar{n}$, and others including $K_{S}^{0}$ and $\Lambda$ that do not yet decay within the tracker volume) and 5 types of confusion (such as different kinds of fake particles, charged particle without track, etc). Each reconstructed particle is characterized by 55 observable variables as inputs to ParT. More details can be found in A.
2.2 Fake particles and BMR
The quantitative analysis [40] shows that fake particles are the leading contribution to the BMR. In our simulation, on average, a $ZH\to\nu\bar{\nu}gg$ event has 80 visible particles with a total energy of 135 GeV at the truth level, while the reconstructed fake particles lead to a double-counted energy of 6 GeV, severely degrading the BMR. PROOF can identify the fake particles with a typical efficiency of 77% and a purity of 97.5%, consequently suppressing the average double-counted energy from fake particles to 0.7 GeV (by nearly one order of magnitude) at the minor cost of mis-excluding visible particles with a total energy on the order of 0.5 GeV. Consequently, PROOF achieves a BMR of 2.75% (see the left panel of Fig. 3), where the effects of fake and mis-excluded particles get balanced. Compared to the CEPC CDR [41], the BMR is improved by 25%, where 10% is from the detector geometry optimization to AURORA [42], and 15% is from the fake particle suppression. The contributions from the three categories of visible-reconstructed particle mapping to the BMR are not proportional to their corresponding energy fractions. The contribution from the 1-1 correspondence category primarily stems from the detector resolution, which is orders of magnitude lower compared to the total energy of particles within this category. Meanwhile, the contributions from the fake and lost categories are comparable to their corresponding total energies. We found that the 1-1 correspondence category contributes to half of the BMR, while the fake and lost categories each contribute a quarter, combined quadratically.
@@ 2.2 Fake particles and BMR
Referring to the truth level energy of visible particles, we calculate the fraction of total visible energy corresponding to different categories of the mapping. Ignoring the remaining fake particles [^2] The remaining fake particles could violate the 1-1 correspondence relationship, as one visible particle could be mapped to one leading reconstructed particle together with one or even several fake particles. However, as the energy of the fake particle is usually much lower than the leading one, the fake particle only has a tiny impact on the reconstruction and identification of the original visible particle. Meanwhile, the total energy of remaining fake particles only contributes to a sub-percentage level of total visible energy. Therefore, ignoring these remaining fake particles in our study., we observe that over 91% of the total visible energy is mapped to well-reconstructed particles, approximately 4% to charged particles without clusters (with over three-fourths of this due to calorimeter acceptance or inefficiency), 2.5% to charged particles without tracks, and 2.2% of the energy is lost due to limited detector acceptance, shower merging, and mis-excluded genuine particles, as illustrated in the pie chart in Fig. 2. The detector acceptance dominates the confusion of charged particles without clusters or tracks. Summing the energy fraction of well-reconstructed particles and confusion type purely originating from detector acceptance, we consider that nearly 95% of the total visible energy maintains the 1-1 correspondence relationship.
2.3 Particle identification
@@
It should be noted that many simplifications have been made to facilitate this analysis. To fully realize 1-1 correspondence reconstruction and maximize its impact on physics exploration, dedicated studies are required in the future. Secondary particles can significantly affect reconstruction and physics measurements. For example, electron-positron pairs from photon conversions and muons from hadron decays could contribute to the background in measurements that rely on (semi-)leptonic decays. Secondary nucleons, generated in the interactions between primary particles and upstream materials, as well as backscattering, could impact the BMR. The reconstruction of the primary vertex is crucial for TOF measurements and affects PID performance. The impact of beam-induced background needs to be analyzed and presumably ameliorated. Event building at high event rates also needs to be developed. Detector optimization should be systematically performed, including the quantification of performance specification (particularly the TOF resolution of the calorimeter and ${\rm d}E/{\rm d}x$ performance of the gaseous tracker), the survey of available and emerging technologies, and the geometry optimization. Modeling and verifying detailed detector responses are crucial for controlling systematic uncertainties, which is vital for precision measurements.
3 Prospects
This section discusses the prospects for high-level reconstruction and physics measurements with an ideal 1-1 correspondence reconstruction.
Since 1-1 correspondence reconstruction provides the particle mass information via PID, the energy of visible particles can be determined by the TOF. 5D calorimetry provides TOF measurements for both charged and neutral particles. For charged particles, their energies/momenta can be determined by three methods: shower energy via calorimeter, track momentum via tracker, and TOF via calorimeter, while the TOF method has a comparative advantage in the forward region. For neutral particles, their energies can be determined by both the shower energy and the TOF via calorimeter, with TOF showing a significant advantage for low-energy neutral hadrons, as demonstrated in the top left panel of Fig. 5. Therefore, by combining these measurements of particle energy, such as selecting the most suitable method according to the particle species and their kinematics, we can significantly improve the energy/momentum resolution of particles, and consequently improve the BMR. Considering the improvement in particle energy measurement can also benefit pattern recognition and confusion control, we estimate that this approach could further enhance the BMR by 15–20%, meaning that a BMR of 2.2–2.4% could, in principle, be achieved.
1-1 correspondence reconstruction can also significantly enhance the JOI performance. JOI is the procedure to determine the type of quark/gluon (11 types are typically considered: $b$, $\bar{b}$, $c$, $\bar{c}$, $s$, $\bar{s}$, $u$, $\bar{u}$, $d$, $\bar{d}$, $g$) from which a jet originates. A recent study realizes JOI with the CEPC CDR baseline detector [38]. It can simultaneously identify $b$, $c$, $s$ quarks with efficiencies of 70–90% and $u$, $d$ quarks with efficiencies of 40%, meanwhile, the charge flip rates of quarks and anti-quarks are controlled to be 10–20%, with an ideal lepton and charged hadron identification. JOI significantly boosts the accuracy of flavor-sensitive physics measurements, such as $H\to s\bar{s}$ and $H\to sb$ decays [38], CKM matrix element $|V_{cb}|$ [43], $CP$-violating phase $\phi_{s}$ [44], and weak mixing angle [45] measurements. 1-1 correspondence provides the identification of $K^{0}_{L}$, $n$, and $\bar{n}$, while in principle the $K^{0}_{S}$ and $\Lambda$ could also be identified [46], therefore, providing the PID information of all these neutral hadrons, we observe a significant improvement on JOI performance especially the identification performance of $u/d/s$ quarks, as shown in the top right panel of Fig. 5.
BMR and JOI are key performances of event reconstruction at electron-positron Higgs factories. Their improvements will benefit most physics measurements with hadronic final states. For example, the Higgs to invisible final states is a key portal to detect dark matter at Higgs factories. The original upper limit at 95% confidence level is quantified to be $\mathcal{O}(0.1\%)$ [47] with a BMR of 3.7%. Improving the BMR to 2.75%, or even 2.2%, will consequently enhance the upper limit by $\mathcal{O}(10\%)$ [48]. Higgs exotic and rare hadronic decays are also sensitive probes to new physics. Using JOI technology, these decay modes can be typically limited to $10^{-4}$ to $10^{-3}$ using $\nu\bar{\nu}H$ and $\ell\ell H$ processes [38]. With the improvement in JOI alone, the constraints could be enhanced by 10% to up to a factor of two. Considering also the improved BMR, the sensitivities are expected to be further enhanced by $\mathcal{O}(10\%)$, see the bottom panel of Fig. 5.
1-1 correspondence can be regarded as a high-efficient information compression that abstracts the entire physics event into $\mathcal{O}(100)$ reconstructed particles, each characterized by approximately $\mathcal{O}(10)$ observables. It enables new analysis methodology to distinguish the signal from background events using all these observables directly. Compared to conventional methodologies, for example the mainstream cut-based and BDT methods that rely on $\mathcal{O}(10)$ of human-defined variables, this new method describes physics events in a much larger parameter space (with the dimension increased by roughly 2 orders of magnitude), thereby significantly enhancing the signal/background separation. A toy analysis demonstrates the effectiveness of this new method in distinguishing the $ZH$ signal from $ZZ$ and $WW$ backgrounds in fully hadronic final states, achieving a performance improvement by more than a factor of 2 [49]. This method can also be applied to detector monitoring and calibration, providing much more information to control the experiment systematic uncertainties. It should be noted that 1-1 correspondence-based analysis relies on accurate MC simulations to provide large statistics of labeled samples. Therefore, the development and validation of suitable MC tools have become critical and should be iteratively refined alongside data analysis, theoretical calculations, and modeling.